# Making Mathematical Reasoning Adaptive

### v1

@@ Abstract
Mathematical reasoning is a primary indicator of large language models (LLMs) intelligence. However, existing LLMs exhibit failures of robustness and generalization. This paper attributes these deficiencies to spurious reasoning—i.e., producing answers from superficial features. To address this challenge, we propose the AdaR framework to enable adaptive reasoning, wherein models rely on problem-solving logic to produce answers. AdaR synthesizes logically equivalent queries by varying variable values, and trains models with RLVR on these data to penalize spurious logic while encouraging adaptive logic. To improve data quality, we extract the problem-solving logic from the original query and generate the corresponding answer by code execution and then apply sanity check. Experimental results demonstrate that AdaR improves robustness and generalization, achieving substantial improvement in mathematical reasoning while maintaining high data efficiency. Analysis indicates that data synthesis and RLVR function in a coordinated manner to enable adaptive reasoning in LLMs. Subsequent analyses derive key design insights into the effect of critical factors and the applicability to instruct LLMs. Our project is available at https://github.com/LaiZhejian/AdaR.
@@ 2.1 Data Synthesis with Executable Code and Verifiable Answers
A straightforward method for perturbing values in variable sets and obtaining correct answers without human annotation is to prompt a LLM. However, our preliminary experiments indicate that this approach offers neither explicit control over the magnitudes and types of perturbations nor any guarantee of answer correctness. To synthesize desired data, the AdaR framework decomposes the task into a sequence of manageable, verifiable sub-tasks as follows.
@@
For certain math word problems, it is crucial to separate the variable set from the solution logic. We regard executable problem-solving code as a problem-solving logic because it generalizes effectively: by substituting the input variable set, it can solve any perturbed query of the original query, for its output can serve as reliable gold answer. Moreover, problem-solving code generation is not challenging when CoT and gold answer are provided, as it reduces to translating the text into code, thereby diminishing the demand on the model’s reasoning capabilities. In addition, the correctness of the generated code can be verified against the gold answer. Building on this insight, we provide a query, the corresponding CoT, and the gold answer to an open-source LLM to synthesize the associated problem-solving code. Meanwhile, because code generation requires abstracting the concrete numerical values in the query into a mapping to the input variables, we further instruct the model to produce a query template as a byproduct by applying this mapping into the original query. Appendix A.2 details the prompt employed.
@@
By comparing the query with the generated template, we construct a variable set that records each variable’s name, numeric type and its value as it appears in the original query. We then apply numerical perturbations, sampling each variable’s value within a range of $\pm\alpha\%$ of its original value. The parameter $\alpha$ allows the control of the perturbation magnitude. Furthermore, to ensure the numerical validity, the numerical type and the sign of each variable before and after perturbation must remain consistent. Formally, for $i$-th original variable’s value $x^{i}_{0}$, we require:
@@
After applying perturbations, we instantiate the template with each variable set to obtain perturbed queries, and execute problem-solving code with the same variable sets as input to generate the corresponding gold answers. To ensure that the perturbed data remain well-posed, we conduct a sanity check along the following aspects:
@@ 5 Related Work
Data synthesis has emerged as a promising solution (Wang et al., 2025b), with Li et al. (2023) paraphrasing questions to diversify templates, though showing limited improvement in adaptive reasoning due to unchanged mathematical structures. Lu et al. (2024) advances this by augmenting datasets through LLM-based generation and verification, helping models identify shared logical patterns through contrasting perturbed queries, but remains constrained by simple perturbations and error-prone verification. More importantly, we put forward a novel perspective, i.e. adaptive reasoning in subfigure I of Figure 1, which can explain why the aforementioned methods are effective. Although RLVR has been widely adopted to enhance model reasoning capabilities (Guo et al., 2025a; Team, 2025; Team et al., 2025) and improve generalization (Wen et al., 2025; Wang et al., 2025a), recent research also reveals several critical limitations, such as format sensitivity (Huang et al., 2025), reward hack (Guo et al., 2025b), and degradation of reasoning capacity boundary (Yue et al., 2025a; Wu et al., 2025). Unlike existing optimization methods that focus primarily on algorithm-level improvements (Yu et al., 2025; Yue et al., 2025b; Zheng et al., 2025), AdaR adopts a complementary approach by inducing adaptive reasoning through the synthesis of high-quality perturbed data.

### v2

@@ Abstract
Mathematical reasoning is a primary indicator of large language models (LLMs) intelligence. However, existing LLMs exhibit failures of robustness and generalization. This paper attributes these deficiencies to spurious reasoning—i.e., producing answers from superficial features. To address this challenge, we propose the AdaR framework to enable adaptive reasoning, wherein models rely on problem-solving logic to produce answers. AdaR synthesizes logically equivalent queries by varying variable values, and trains models with RLVR on these data to penalize spurious logic while encouraging adaptive logic. To improve data quality, we extract the problem-solving logic from the original query and generate the corresponding answer by code execution and then apply sanity check. Experimental results demonstrate that AdaR improves robustness and generalization, achieving substantial improvement in mathematical reasoning while maintaining high data efficiency. Analysis indicates that data synthesis and RLVR function in a coordinated manner to enable adaptive reasoning in LLMs. Subsequent analyses derive key design insights into the effect of critical factors and the applicability to instruct LLMs. Our project is available at https://github.com/NJUNLP/AdaR.
@@ 2.1 Data Synthesis with Executable Code and Verifiable Answers
A straightforward method for perturbing values in queries and obtaining correct answers without human annotation is to prompt a LLM (Wang et al., 2025b). However, our preliminary experiments indicate that this approach offers neither explicit control over the magnitudes and types of perturbations nor any guarantee of answer correctness. To synthesize the desired query-answer pairs, AdaR constrains the model to identify only the metadata within the data and to translate a CoT into code; controllable perturbations are then applied externally, and code execution is used to ensure answer correctness. The AdaR framework accordingly decomposes it into a sequence of manageable, verifiable sub-tasks as follows.
@@
Executable problem-solving code can serve as a problem-solving logic provided its correctness is guaranteed. Crucially, such code generalizes effectively: by substituting the input variable set, it can solve any perturbed query of the original query, for its output can serve as reliable gold answer. When CoT and gold answer are provided, problem-solving code generation becomes a straightforward process of translating the textual logic (i.e. CoT) into executable code, thereby reducing the model’s reasoning burden and enabling direct verification of correctness against the gold answer. Building on these insights, we provide a query, the corresponding CoT and gold answer to an open-source LLM to synthesize the associated problem-solving code for producing gold answers of subsequent perturbed queries. Because code generation requires abstracting the concrete numerical values in the query into a mapping to the input variables, we further instruct the model to produce a query template as a byproduct by applying this mapping into the original query. Appendix A.2 details the prompt employed.
@@
By comparing the query with the generated template, we construct a variable set that records each variable’s name, numeric type and its value as it appears in the original query. This variable set, enriched with metadata, enables direct numeric perturbations at a predefined magnitude while ensuring numerical validity, thereby avoiding the unpredictability of perturbations performed through an LLM. Specifically, we then apply independent perturbations by sampling each variable’s value within a range of $\pm\alpha\%$ of its original value. The parameter $\alpha$ allows the control of the perturbation magnitude. To ensure numerical validity, both the numeric type and the sign of every variable must remain unchanged before and after perturbation. Formally, for the $i$-th variable with original value $x_{i}^{(0)}$, we draw:
@@
After applying perturbations, we instantiate the template with each variable set to obtain perturbed queries, and execute problem-solving code with the same variable sets as input to generate the corresponding gold answers. Although providing the CoT and answer has reduced task difficulty, query template and problem-solving code generated by the LLM remain uncertain and may contain errors. Furthermore, a meaningful query imposes inter-variable constraints on its variables; independent perturbations can violate these constraints and thereby introduce errors. To ensure that the perturbed data remain well-posed, we conduct a sanity check along the following aspects:
@@ 5 Related Work
Data synthesis has emerged as a promising solution (Wang et al., 2025b; Geng et al., 2025), with Li et al. (2023) paraphrasing questions to diversify templates, though showing limited improvement in adaptive reasoning due to unchanged mathematical structures. Lu et al. (2024) advances this by augmenting datasets through LLM-based generation and verification, helping models identify shared logical patterns through contrasting perturbed queries, but remains constrained by simple perturbations and error-prone verification. More importantly, we put forward a novel perspective, i.e. adaptive reasoning in subfigure I of Figure 1, which can explain why the aforementioned methods are effective. Although RLVR has been widely adopted to enhance model reasoning capabilities (Guo et al., 2025a; Team, 2025; Team et al., 2025a; b) and improve generalization (Wen et al., 2025; Wang et al., 2025a), recent research also reveals several critical limitations, such as format sensitivity (Huang et al., 2025), reward hack (Guo et al., 2025b), and degradation of reasoning capacity boundary (Yue et al., 2025a; Wu et al., 2025). Unlike existing optimization methods that focus primarily on algorithm-level improvements (Yu et al., 2025; Yue et al., 2025b; Zheng et al., 2025), AdaR adopts a complementary approach by inducing adaptive reasoning through the synthesis of high-quality perturbed data.